# The WMDP Benchmark: Measuring and Reducing Malicious Use With Unlearning

### v2

@@ Abstract
The White House Executive Order on Artificial Intelligence highlights the risks of large language models (LLMs) empowering malicious actors in developing biological, cyber, and chemical weapons. To measure these risks, government institutions and major AI labs are developing evaluations for hazardous capabilities in LLMs. However, current evaluations are private and restricted to a narrow range of malicious use scenarios, limiting further research into mitigating risk. To fill these gaps, we publicly release the Weapons of Mass Destruction Proxy (WMDP) benchmark, a dataset of multiple-choice questions that serve as a proxy measurement of hazardous knowledge in biosecurity, cybersecurity, and chemical security. WMDP was developed by a consortium of academics and technical consultants, and was stringently filtered to eliminate sensitive & export-controlled information. WMDP serves two roles: first, as an evaluation for hazardous knowledge in LLMs, and second, as a benchmark for unlearning methods to remove such hazardous knowledge. To guide progress on unlearning, we develop Cut, a state-of-the-art unlearning method based on controlling model representations. Cut reduces model performance on WMDP while maintaining general capabilities in areas such as biology and computer science, suggesting that unlearning may be a concrete path towards reducing malicious use from LLMs. We release our benchmark and code publicly at https://wmdp.ai.
@@ 1 Introduction
Similar to other technologies, such as gene editing and nuclear energy, AI is dual-use—it can be leveraged for benefit and harm (Urbina et al., 2022). To address its dual-use risks, the White House Executive Order on Artificial Intelligence (White House, 2023) calls for investigation into the ability of AI to enable malicious actors in developing chemical, biological, radiological, nuclear, and cyber weapons. For instance, AI coding assistants may lower the barrier of entry for novices to conduct cyberattacks (Fang et al., 2024), potentially increasing attack scale and frequency, and risking catastrophe if directed towards critical infrastructure, such as power grids (UK Cabinet Office, 2023). Likewise, AI assistants for biology could troubleshoot bottlenecks in biological weapons development, increasing the frequency of attempts to build a bioweapon and straining risk mitigation measures (Sandbrink, 2023). This has motivated government institutions and major AI labs to anticipate risk by designing evaluations for AI-aided biological threats (UK AI Safety Summit, 2023; Anthropic, 2023; OpenAI, 2023b, 2024; Mouton et al., 2024).
Unfortunately, many existing evaluations of hazardous capabilities do not provide a guide for risk mitigation. Currently, developers use private, manual, and highly-specific risk evaluations, investigating whether models can build biological weapons end-to-end (Sandbrink, 2023) or hack well enough to exfiltrate their own weights (Shevlane et al., 2023). However, private benchmarking limits scientific inquiry into measuring and reducing malicious use. Furthermore, these evaluations test a small number of specific risk pathways, so low performance does not guarantee that LLMs are secure across the broad distribution of malicious use risks.
Additionally, technical solutions to reduce malicious use in LLMs are limited. The primary safeguard is training models to refuse harmful queries (Ouyang et al., 2022; Bai et al., 2022; Mazeika et al., 2024), but adversaries can deploy adversarial attacks (Wei et al., 2023; Zou et al., 2023b) to bypass models’ refusal training. Another proposal is to filter hazardous information from the pretraining data (Ngo et al., 2021), but adversaries may reintroduce this information through finetuning (Zhan et al., 2023; Qi et al., 2023; Pelrine et al., 2023). Fortunately, closed-source LLM providers have leverage, as they may apply safety interventions before serving LLMs. In particular, providers may perform unlearning to directly remove hazardous knowledge (Figure 2). Unlearned models have higher inherent safety: even if they are jailbroken, unlearned models lack the hazardous knowledge necessary to enable malicious users (Hendrycks et al., 2021). However, research into unlearning hazardous knowledge is bottlenecked by the lack of a public benchmark.
@@
To guide progress on unlearning, we develop Contrastive Unlearn Tuning (Cut), a state-of-the-art unlearning method that removes hazardous knowledge while preserving general model capabilities. Inspired by representation engineering (Zou et al., 2023a), Cut pushes model activations along an unlearning direction on hazardous data, while regularizing towards original, frozen model weights on benign data (Section 4). Cut significantly reduces model performance on WMDP, while mostly retaining general capabilities on MMLU (Hendrycks et al., 2020b) and MT-Bench (Zheng et al., 2023a), suggesting that concrete progress on unlearning is feasible (Section 5.2) with minimal impact on the general scientific capabilities of LLMs. We provide evidence that Cut is robust, as unlearned knowledge cannot be recovered by linear probing or adversarial attacks (Sections 5.2 and 5.3).
Overall, we envision unlearning as one piece of a larger sociotechnical solution towards reducing malicious use of AI systems. Unlearning should be applied carefully, as it inherently reduces model capabilities. Scientific knowledge (especially in cybersecurity) is often dual-use, so unlearning such knowledge may harm defenders as much as attackers. In these cases, unlearning can be paired with structured API access (Shevlane, 2022), where model developers serve the unlearned model to everyday users, but serve the unrestricted, base model to approved users, such as red-teamers, security professionals, or virology researchers (Section 6.2). Furthermore, other risk mitigation strategies may be promoted, such as implementing universal screening of synthetic DNA orders to prevent widespread access to pathogen DNA. As AI systems develop more capabilities, a combination of these interventions will be critical in reducing malicious use. To enable further research on unlearning hazardous knowledge, we release our datasets, code, and models publicly at https://wmdp.ai.
@@ 2 Related Work
Specifically, one growing concern is the ability of LLMs to assist with malicious use. In particular, LLMs may aid actors in planning bioattacks (Sandbrink, 2023) and procuring pathogens (Gopal et al., 2023). Moreover, LLMs can assist users in synthesizing dangerous chemicals (Boiko et al., 2023) or conducting cyberattacks (Bhatt et al., 2023). In response to these emergent hazardous capabilities (Hendrycks et al., 2021), major AI labs have developed frameworks to measure and mitigate biological, cybersecurity, and chemical hazards posed by their models (Anthropic, 2023; OpenAI, 2023b, 2024). Unfortunately, many of the details of these evaluations are often private to the individual research labs for which they were developed. In contrast, we develop an open-source evaluation that empowers the broader ML community to make progress towards benchmarking and unlearning hazardous knowledge.
@@
Unlearning (Cao and Yang, 2015) originally gained traction as a response to privacy concerns in light of regulation (Council of European Union, 2014; CCPA, 2018), and most methods focused on erasing specific samples or facts (Golatkar et al., 2020; Liu et al., 2020; Meng et al., 2022; Jang et al., 2023; Pawelczyk et al., 2023) rather than entire domains. Goel et al. (2024) show existing unlearning methods struggle to remove knowledge without access to all relevant training data, a challenge Cut overcomes.
@@ 3 The WMDP Benchmark
WMDP is an automatic, public benchmark of hazardous capabilities that serves as a guide for risk mitigation (Section 3.1). We create questions by designing threat models for biosecurity.(Section 3.2), cybersecurity (Section 3.3), and chemistry (Section 3.4). We also remove sensitive and export-controlled information from entering WMDP (Section 3.5). To further unlearning research beyond WMDP, we also provide additional unlearning benchmarks based on MMLU (Appendix C).
@@ 3.2 Biosecurity Threat Model
In biosecurity, the malicious use threats that are increased by AI can be broadly categorized as expanding access to pre-existing threats (by lowering barriers to entry), and unlocking new areas of biology (by synthesizing new knowledge or accelerating in-silico modeling and experimentation).
We primarily focus on the development and dissemination of transmissible potential pandemic agents, such as influenza, smallpox, etc. While our dataset additionally includes some information about highly lethal non-transmissible bioweapons like anthrax, we believe the majority of emerging risk from biotechnology stems from advances in synthetic biology and bioengineering that increase access to, or modify, the design and development of transmissible agents.
A standard biotechnology risk chain can be seen in Figure 5. In this threat model, “ideation” involves actively planning for a biological attack; “design” involves retrieving blueprints for a hazardous agent, such as determining the DNA sequence; “build” consists of the protocols, reagents, and equipment necessary to create the threat; and “test” consists of measuring characteristics or properties of the pathogen of interest. By “learning” from these results and iterating after the test phase, the design can be revised until the threat is released.
@@ 3.5 Sensitive Information Mitigation
4 Cut: Unlearning Inspired By Representation Engineering
We introduce Contrastive Unlearn Tuning (Cut), a finetuning method for unlearning hazardous knowledge (Algorithm 1). We outline the setup (Section 4.1) and explain our method (Section 4.2). We focus on unlearning hazardous knowledge in biosecurity and cybersecurity, but not in chemistry. While WMDP-Chem is a useful tool for hazard measurement, we are more uncertain if the hazard mitigation benefits of unlearning on WMDP-Chem outweigh the costs on general model capabilities.
@@ 4.2 Method
We leverage the idea that model representations encode knowledge of the world and that these representations may be controlled to guide the model (Ilharco et al., 2023; Zou et al., 2023a; Turner et al., 2023). Intuitively, we control the model towards having a novice-like level of hazardous knowledge. We design a loss function with a forget loss and a retain loss. The forget loss bends the model representations towards those of a novice, while the retain loss limits the amount of general capabilities removed (Figure 7).
@@
Our goal is to bend the model representations on hazardous knowledge towards those of a novice. We must precisely specify both the distribution of knowledge to unlearn and the direction to push the activations towards. If we misspecify the distribution, we will either delete too little or too much information. If we misspecify the direction, we may degrade the model in unintended ways.
To specify the distribution of knowledge to unlearn (e.g., cybersecurity), we use keywords (e.g., exploit development or penetration testing). If we remove knowledge specified by such keywords, then we can maximally remove hazardous knowledge while minimally destroying general capabilities.
To guide activations, we use control vectors (Zou et al., 2023a; Ilharco et al., 2023). Given some keyword (e.g., “penetration testing”), an unlearning control vector is a vector in the model’s activation space that represents the absence of knowledge of that keyword. Controlling the model with unlearning control vectors is more precise than controlling towards a generic direction.
Inspired by representation engineering (Zou et al., 2023a), our forget loss first calculates unlearning control vectors based on a set of keywords, and then controls the model activations along the direction of the unlearning control vectors. To calculate the unlearning control vectors, we assume access to , the hidden states of the unlearned model at some layer and , the hidden states of the original, frozen model at some layer . Given a , we compute , , and
Our keywords are generated by the topics in threat models for biosecurity and cybersecurity (Appendix B.1). For example, for biosecurity, one of our keywords is “bioterrorism” and for cybersecurity, one of our keywords is “penetration testing.”
To control the activations, we assume a forget dataset and compute:
where is a randomly sampled unlearning control vector from our set.
@@
Cut finetunes the model weights to minimize this loss. To unlearn multiple distributions of knowledge, we interleave the gradient updates (i.e., update model weights on the biosecurity distribution, then update on the cybersecurity distribution, then repeat). In practice, we find it sufficient to compute the loss only on layer and update gradients only on layers , , and . We leverage this observation to save memory and efficiently unlearn on large (34B parameter) models.
@@
Similarly, to preserve activations on general language modelling tasks, we need to collect , a knowledge preservation distribution which approximates general, non-hazardous knowledge. For these, we collected subject-specific retain sets detailed in Sections A.4 and A.5. However, we find in practice that Cut is more performant when has qualitatively distinct content from , so as not to relearn the unlearned knowledge. Thus, we set to be Wikitext (Merity et al., 2016). We release the unused subject-specific retain sets for WMDP-Bio and WMDP-Cyber publicly, to guide future unlearning methods that can more effectively use these corpora.
@@ 5 Experimental Results
We examine the performance of Cut and other unlearning methods. We describe the experimental setup (Section 5.1) and provide quantitative (Section 5.2) and robustness (Section 5.3) evaluations. We also check if unlearning on WMDP generalizes to more hazardous information (Section 5.4). Cut markedly improves upon existing baselines, but future work is necessary to improve the precision of unlearning hazardous knowledge while fully maintaining general capabilities.
@@ 5.1 Setup
We describe the benchmarks we use for evaluations, the models we use for unlearning, and the baselines we use for comparisons. We only conduct unlearning experiments on WMDP-Bio and WMDP-Cyber, as discussed in Section 4.Benchmarks.
We evaluate removal of hazardous knowledge with WMDP. To evaluate the preservation of general knowledge, we use MMLU (Hendrycks et al., 2020b), focusing on topics similar to biosecurity (college biology, virology) and cybersecurity (college computer science, computer security). Finally, to evaluate the fluency of models, we use MT-Bench, a multi-turn conservation and instruction-following benchmark (Zheng et al., 2023b).Models.
We remove knowledge of biosecurity and cybersecurity on zephyr-7b-beta (Tunstall et al., 2023) and Yi-34b-Chat (01-ai, 2023), two of the most performant open-source generative language models at 7 billion, and 34 billion, parameters respectively. Additionally, we report the performance of GPT-4 (OpenAI, 2023a) as an upper bound on benchmark performance.Baselines.
We benchmark Cut against three unlearning baselines: SCRUB (Kurmanji et al., 2023), SSD (Foster et al., 2024), and LLMU (Yao et al., 2023b), on zephyr-7b. Because we found low performance on zephyr-7b, we did not benchmark the baselines on Yi-34b. See Appendix B.6 for our implementation of the baselines.

### v3

@@ Abstract
The White House Executive Order on Artificial Intelligence highlights the risks of large language models (LLMs) empowering malicious actors in developing biological, cyber, and chemical weapons. To measure these risks, government institutions and major AI labs are developing evaluations for hazardous capabilities in LLMs. However, current evaluations are private and restricted to a narrow range of malicious use scenarios, limiting further research into mitigating risk. To fill these gaps, we publicly release the Weapons of Mass Destruction Proxy (WMDP) benchmark, a dataset of multiple-choice questions that serve as a proxy measurement of hazardous knowledge in biosecurity, cybersecurity, and chemical security. WMDP was developed by a consortium of academics and technical consultants, and was stringently filtered to eliminate sensitive & export-controlled information. WMDP serves two roles: first, as an evaluation for hazardous knowledge in LLMs, and second, as a benchmark for unlearning methods to remove such hazardous knowledge. To guide progress on unlearning, we develop RMU, a state-of-the-art unlearning method based on controlling model representations. RMU reduces model performance on WMDP while maintaining general capabilities in areas such as biology and computer science, suggesting that unlearning may be a concrete path towards reducing malicious use from LLMs. We release our benchmark and code publicly at https://wmdp.ai.
@@ 1 Introduction
Similar to other technologies, such as gene editing and nuclear energy, AI is dual-use—it can be leveraged for benefit and harm (Urbina et al., 2022). To address its dual-use risks, the White House Executive Order on Artificial Intelligence (White House, 2023) calls for investigation into the ability of AI to enable malicious actors in developing chemical, biological, radiological, nuclear, and cyber weapons. For instance, AI coding assistants may lower the barrier of entry for novices to conduct cyberattacks (Fang et al., 2024), potentially increasing the frequency of cyberattacks and the risk of catastrophe, especially if these attacks directed towards critical infrastructure, such as power grids (UK Cabinet Office, 2023). Likewise, AI assistants for biology could troubleshoot bottlenecks in biological weapons development, increasing the frequency of attempts to build a bioweapon and straining risk mitigation measures (Sandbrink, 2023). This has motivated government institutions and major AI labs to anticipate risk by designing evaluations for AI-aided biological threats (UK AI Safety Summit, 2023; Anthropic, 2023; OpenAI, 2024; Mouton et al., 2024; Phuong et al., 2024).
Unfortunately, current evaluations of hazardous capabilities do not provide a guide for mitigating malicious use risk. For example, developers evaluate whether models can build biological weapons end-to-end (Sandbrink, 2023) or hack well enough to exfiltrate their own weights (Shevlane et al., 2023), creating private, manual, and highly-specific evaluations. Because these evaluations test a small number of specific risk pathways, low performance on them does not guarantee that LLMs are secure across the broad distribution of malicious use risks. More importantly, such private benchmarking limits scientific inquiry towards measuring and reducing malicious use.
Developers also lack robust technical solutions to reduce malicious use in LLMs. The primary safeguard is training models to refuse harmful queries (Ouyang et al., 2022; Bai et al., 2022; Mazeika et al., 2024), but adversaries can deploy adversarial attacks (Wei et al., 2023; Zou et al., 2023b) to bypass models’ refusal training. Another proposal is to filter hazardous information from the pretraining data (Ngo et al., 2021), but adversaries may reintroduce this information through finetuning (Zhan et al., 2023; Qi et al., 2023; Pelrine et al., 2023). A promising approach for closed-source LLM providers is unlearning, directly removing hazardous knowledge before model serving (Figure 2). Unlearned models have higher inherent safety: even if they are jailbroken, unlearned models lack the hazardous knowledge necessary to enable malicious users (Hendrycks et al., 2021). However, research into unlearning hazardous knowledge is bottlenecked by the lack of a public benchmark.
@@
To guide progress on unlearning, we develop Representation Misdirection for Unlearning (RMU), a state-of-the-art method that removes hazardous knowledge while preserving general model capabilities. Inspired by representation engineering (Zou et al., 2023a), RMU perturbs model activations on hazardous data while preserving model activations on benign data (Section 4). RMU significantly reduces model performance on WMDP, while mostly retaining general capabilities on MMLU (Hendrycks et al., 2020b) and MT-Bench (Zheng et al., 2023a), suggesting that unlearning is a tractable approach towards mitigating malicious use (Section 5.2). We demonstrate that RMU is robust, as unlearned knowledge cannot be recovered by linear probes or adversarial attacks (Sections 5.2 and 5.3).
Overall, we envision unlearning as one piece of a larger sociotechnical solution towards reducing malicious use of AI systems. Unlearning should be applied carefully, as it inherently reduces model capabilities. Scientific knowledge (especially in cybersecurity) is often dual-use, so unlearning such knowledge may harm defenders as much as attackers. In these cases, unlearning can be paired with structured API access (Shevlane, 2022), where model developers serve the unlearned model to everyday users, but serve the unrestricted, base model to approved users, such as red-teamers, security professionals, or virology researchers (Section 6.2). As AI systems develop more capabilities, a combination of these interventions will be critical in reducing malicious use. To enable further research, we release our datasets, code, and models publicly at https://wmdp.ai.
@@ 2 Related Work
Specifically, one growing concern is the ability of LLMs to assist with malicious use. In particular, LLMs may aid actors in planning bioattacks (Sandbrink, 2023) and procuring pathogens (Gopal et al., 2023). Moreover, LLMs can assist users in synthesizing dangerous chemicals (Boiko et al., 2023) or conducting cyberattacks (Bhatt et al., 2023). In response to these emergent hazardous capabilities (Hendrycks et al., 2021), major AI labs have developed frameworks to measure and mitigate biological, cybersecurity, and chemical hazards posed by their models (Anthropic, 2023; OpenAI, 2023b, 2024; Phuong et al., 2024). Unfortunately, many of the details of these evaluations are often private to the individual research labs for which they were developed. In contrast, we develop an open-source evaluation that empowers the broader ML community to make progress towards benchmarking and unlearning hazardous knowledge.
@@
Unlearning (Cao and Yang, 2015) originally gained traction as a response to privacy concerns in light of regulation (Council of European Union, 2014; CCPA, 2018), and most methods focused on erasing specific samples or facts (Golatkar et al., 2020; Liu et al., 2020; Meng et al., 2022; Jang et al., 2023; Pawelczyk et al., 2023) rather than entire domains. Goel et al. (2024) show existing unlearning methods struggle to remove knowledge without access to all relevant training data, a challenge RMU overcomes.
@@ 3 The WMDP Benchmark
WMDP is an automatic, public benchmark of hazardous capabilities that serves as a guide for risk mitigation (Section 3.1). We create questions by designing threat models for biosecurity (Section 3.2), cybersecurity (Section 3.3), and chemistry (Section 3.4). We also remove sensitive and export-controlled information from entering WMDP (Section 3.5). To further unlearning research beyond WMDP, we also provide additional unlearning benchmarks based on MMLU (Appendix C).
@@ 3.2 Biosecurity Threat Model
In biosecurity, the malicious use threats that are increased by AI can be broadly categorized as expanding access to pre-existing threats (by lowering barriers to entry), and unlocking new areas of biology (by synthesizing new knowledge or accelerating in-silico modeling and experimentation) (Sandbrink, 2023).
We primarily focus on the development and dissemination of transmissible potential pandemic agents, such as influenza, smallpox, etc. While our dataset additionally includes some information about highly lethal non-transmissible bioweapons like anthrax, we believe the majority of emerging risk from biotechnology stems from advances in synthetic biology and bioengineering that increase access to, or modify, the design and development of transmissible agents (Esvelt, 2022).
A standard biotechnology risk chain can be seen in Figure 5. In this threat model, “ideation” involves actively planning for a biological attack; “design” involves retrieving blueprints for a hazardous agent, such as determining the DNA sequence; “build” consists of the protocols, reagents, and equipment necessary to create the threat; and “test” consists of measuring characteristics or properties of the pathogen of interest. By “learning” from these results and iterating after the test phase, the design can be revised until the threat is released (Nelson and Rose, 2023).
@@ 3.5 Sensitive Information Mitigation
4 RMU: Unlearning Inspired By Representation Engineering
We introduce Representation Misdirection for Unlearning (RMU), a finetuning method for unlearning hazardous knowledge (Algorithm 1). We outline the setup (Section 4.1) and explain our method (Section 4.2), with further detail in Appendix B.4 and B.5. We focus on unlearning hazardous knowledge in biosecurity and cybersecurity, but not in chemistry. While WMDP-Chem is a useful tool for hazard measurement, we are more uncertain if the hazard mitigation benefits of unlearning on WMDP-Chem outweigh the costs on general model capabilities.
@@ 4.2 Method
Classically, language models are trained with a loss on their outputs (Vaswani et al., 2017; Devlin et al., 2018). On the other hand, mechanistic interpretability proposes editing models by intervening on individual neurons (Wang et al., 2022). In contrast to both these perspectives, we leverage the idea that model representations encode knowledge of the world (Meng et al., 2022) and that these representations may be manipulated to affect model behavior (Zou et al., 2023a; Ilharco et al., 2023; Turner et al., 2023). We design a two-part loss function with a forget loss and a retain loss; intuitively, the forget loss perturbs the model activations on hazardous data while the retain loss preserves its activations on benign data (Figure 7).
@@
Our goal is to degrade the model’s representations of hazardous knowledge. Our experiments suggest that increasing the norm of the model’s activations on hazardous data in earlier layers makes it difficult for later layers to process the activations, achieving our desiderata.
To calculate our forget loss, we assume access to , the hidden states of the unlearned model at some layer and , the hidden states of the original, frozen model at some layer . Then, we compute , a random unit vector with independent entries sampled uniformly at random from . Note that is held fixed throughout training. Given a forget dataset , we compute:
where is the number of tokens in and is some hyperparameter that controls activation scaling.
@@
RMU finetunes the model weights to minimize this loss. To unlearn multiple distributions of knowledge, we interleave the gradient updates (i.e., update model weights on the biosecurity distribution, then update on the cybersecurity distribution, then repeat). In practice, we find it sufficient to compute the loss only on layer and update gradients only on layers , , and . We leverage this observation to save memory and efficiently unlearn on larger LMs.
@@
Similarly, to preserve activations on general language modelling tasks, we need to collect , a knowledge preservation distribution which approximates general, non-hazardous knowledge. For these, we collected subject-specific retain sets detailed in Sections A.4 and A.5. However, we find in practice that RMU is more performant when has qualitatively distinct content from , so as not to relearn the unlearned knowledge. Thus, we set to be Wikitext (Merity et al., 2016). We release the unused subject-specific retain sets for WMDP-Bio and WMDP-Cyber publicly, to guide future unlearning methods that can more effectively use these corpora.
@@ 5 Experimental Results
We examine the performance of RMU and other unlearning methods. We describe the experimental setup (Section 5.1) and provide quantitative (Section 5.2) and robustness (Section 5.3) evaluations. We also check if unlearning on WMDP generalizes to more hazardous information (Section 5.4). Finally, we report plots for how RMU scales activations on hazardous and benign data in Appendix B.5. RMU markedly improves upon existing baselines, but future work is necessary to improve the precision of unlearning hazardous knowledge while fully maintaining general capabilities.
| Model | WMDP () | MMLU () | MT-Bench () | |
| Bio | Cyber | |||
| zephyr-7b | ||||
[…]
| SCRUB | ||||
| SSD | ||||
| RMU (ours) | ||||
| Yi-34b | ||||
[…]
| Mixtral-8x7B | ||||
[…]
@@ 5.1 Setup
We describe the benchmarks we use for evaluations, the models we use for unlearning, and the baselines we use for comparisons. We only conduct unlearning experiments on WMDP-Bio and WMDP-Cyber, as discussed in Section 4.
Benchmarks.
We evaluate removal of hazardous knowledge with WMDP. To evaluate the preservation of general knowledge, we use MMLU (Hendrycks et al., 2020b), focusing on topics similar to biosecurity (college biology, virology) and cybersecurity (college computer science, computer security). Finally, to evaluate the fluency of models, we use MT-Bench, a multi-turn conservation and instruction-following benchmark (Zheng et al., 2023b).
Models.
We remove knowledge of biosecurity and cybersecurity on zephyr-7b-beta (Tunstall et al., 2023), Yi-34b-Chat (01-ai, 2023), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024), three of the most performant open-source generative language models at their respective sizes. Additionally, we report the performance of GPT-4 (OpenAI, 2023a) as an upper bound on benchmark performance.
Baselines.
We benchmark RMU against three unlearning baselines: SCRUB (Kurmanji et al., 2023), SSD (Foster et al., 2024), and LLMU (Yao et al., 2023b), on zephyr-7b. Because we found low performance on zephyr-7b, we did not benchmark the baselines on Yi-34b or Mixtral-8x7B. See Appendix B.7 for our implementation of the baselines.